OpenStudy (anonymous):

Find the Horizontal Asymptote of x^2+x-2 divided by x^2-3x-4.

OpenStudy (amistre64):

umm, we (err i) did that in the last post

OpenStudy (anonymous):

The last one was finding the range. this is the horizontal asymptote.

OpenStudy (amistre64):

i found the HA during the process to test if it was within the range

OpenStudy (anonymous):

Mkaaay, I will look to that one for reference then. I'm sorry!

OpenStudy (anonymous):

There was so much going on on the last one. hehe :P

OpenStudy (amistre64):

the idea is that when x is massively huge; the function is controlled top and bottom by the highest exponentail term

OpenStudy (amistre64):

so; at large values of x \[ \frac{x^2+x-2}{x^2-3x-4}=\frac{x^2}{x^2}\] dont know a latex for "approximately

OpenStudy (anonymous):

Okay.

OpenStudy (amistre64):

this simplifies to 1

OpenStudy (anonymous):

Oh, simple.

OpenStudy (anonymous):

So since numerator degree=denominator degree. It's the ratio of coefficients?

OpenStudy (amistre64):

there are 3 conditions that can occur in rational expressions; the one we did was equal degrees if the degrees are unequal, say; high top, low bottom\[\frac{x^6}{x^2}=x^4\]the top wins out and this thing goes up to infinity if low top, high bottom\[\frac{x^3}{x^{10}}=\frac1{x^7}\]the bottom wins out and this thing goes to zero

OpenStudy (amistre64):

yes, top equals bottom, it ration of first coeffs

OpenStudy (anonymous):

Ohhh, okay. :)